OpenStudy (anonymous):

sqrt 3 times sqrt 21

OpenStudy (driftracer305):

just multiply them together.........

OpenStudy (anonymous):

You can rewrite the expression this way: \[\sqrt{3}\sqrt{21}=\sqrt{3}\sqrt{3}\sqrt{7}\] Then, because the square root of 3 times itself is 3, you get this: \[\sqrt{3}\sqrt{3}\sqrt{7}=3\sqrt{7}\]

OpenStudy (driftracer305):

but u knw u will have 2 sqrt 3's..........coz of sqrt 7 * 3 = 21....so u will have one 3 outside as well

OpenStudy (anonymous):

oh. well. thats confusing aha but i think i get it ... thank you! @jabberwock @driftracer305

OpenStudy (driftracer305):

yea....jabberwock is right........

OpenStudy (anonymous):

No worries :)